OpenStudy (lilkg77):

PLEASE HELP!!!! Which of the following equations describe the line shown below? CHECK ALL THAT APPLY. Line =(-3, 2) (1, 10) A. y-3=2(x-2) B. y-2=2(x-3) C. y-10=2(x-1) D. y-2=2(x+3) E. y-1=2(x-10)

jimthompson5910 (jim_thompson5910):

Checking Choice A: y - 3 = 2(x-2) Let's check the first point (-3,2) So x = -3 and y = 2. Plug these values into the equation y - 3 = 2(x-2) 2 - 3 = 2(-3-2) -1 = 2(-5) -1 = -10 The last equation -1 = -10 is false. So y - 3 = 2(x-2) is false when (x,y) = (-3,2). This means (-3,2) is NOT on the line y - 3 = 2(x-2) So it's impossible for choice A to be the answer. We can eliminate choice A Go through the other choices and check both points. Make sure that both points make the equation true. If not, then you can eliminate the answer choice like I did above.

OpenStudy (anonymous):

http://www.cymath.com/answer.php?q=(-3%2C%202)%20(1%2C%2010) This Might Help